The AEB frame is affected by a vertical force F1=47 kN and

a horizontal force F2=28 kN, a distributed load

max q=28N/m. Pic.4

Determine the reaction of the support A and B, if the distance AC=CD=19 m, DE=38, the point K, to which the force F2 is applied, is

the middle of the segment EB, and the angle between AE and the line AB is

30. Consider the triangle AEB rectangular..


1
Expert's answer
2021-06-07T09:33:59-0400

"F_A=F_1-F_2+q\frac{DE^2}{2AC},"

"F_B=F_2-F_1+q\frac{DE^2}{2CD},"

"F_A=47-28+28\cdot \frac{38^2}{2\cdot 19}=1083~N,"

"F_B=28-47+28\cdot \frac{38^2}{2\cdot 19}=1045~N."